# Lumina-mGPT: Illuminate Flexible Photorealistic Text-to-Image Generation with Multimodal Generative Pretraining

### v1

@@ 3.2 Lumina-mGPT
Building on pretrained Multimodal Generative Pretraining (mGPT) representation, Lumina-mGPT enables flexible photoelastic image generation as well as vision and language tasks unification. To fully unlock the potential of mGPT, we propose Unambiguous Image Representation (Uni-Rep), Flexible Progressive Supervised Finetuning (FP-SFT) and Omnipotent Supervised Finetuning (Omni-SFT). Uni-Rep eliminates the ambiguity in the original 2D shape of the flattened 1D image tokens, building the basis for image understanding and generation at flexible resolution and aspect ratio. FP-SFT progressively finetunes mGPT to generate higher-resolution discrete image tokens in a simple-to-hard manner, while incorporating multi-task finetuning to prevent the catastrophic forgetting of text knowledge. After the FP-SFT process, Lumina-mGPT becomes capable of generating photorealistic images at flexible resolutions. Building on the image generation ability from the FP-SFT stage, we continue to finetune Lumina-mGPT over omnipotent task datasets by tokenizing images and annotations from dense labeling, spatial-conditional image generation, and multi-turn editing datasets into discrete tokens. All finetuning stages adopt the next-token prediction objective to jointly model multimodal tokens and integrate z-loss, similar to the pretraining stage.
@@ 3.2.1 Effective Initlization
Large-scale pre-training and scalable model architecture have been widely verified as the golden path to advanced intelligence. As mGPTs, like Chameleon (Team, 2024a), are pretrained on large-scale image-text interleaved datasets and have developed effective and generalizable representations for both image and text, they can better serve as the starting pointing for flexible photorealistic image generation and beyond than random initialization or language-only models. Furthermore, the LLaMA (Touvron et al., 2023) architecture, which incorporates features like query-key normalization (Henry et al., 2020) and rotary position encoding (Su et al., 2024), has demonstrated strength and scalability through extensive validation (Touvron et al., 2023; Meta, 2024; Gao et al., 2024). By initializing from the Chameleon mGPT, which adheres to the LLaMA architecture, we can leverage these architectural advantages. Consequently, initializing from mGPTs allowed us to efficiently train high-performing Lumina-mGPT models, with parameters ranging from 7B to 30B, using just 10M high-quality image-text data points.
We emphasize that pre-training a decoder-only multimodal Generative PreTraining (mGPT) model has been extensively explored and applied across various applications (Team, 2024a; Wang et al., 2022; Lu et al., 2024a). Unlike these existing approaches, which aim to unify different modalities and tasks, the primary contribution of Lumina-mGPT lies in demonstrating the benefits of starting with an appropriate mGPT representation rather than using large language models (LLMs) or random initialization, especially in text-to-image generation. We believe this finding can inspire future advancements in image generation and vision generalist models.
@@ 3.2.2 Supervised Finetuning for Lumina-mGPT
Existing methods, such as Chameleon (Team, 2024a) and LlamaGen (Sun et al., 2024), represent images as 1D flattened sequences of 2D discrete image codes. While this approach is adequate for fixed image resolution, it becomes ambiguous when the image resolution is supported to be variable, as with Lumina-mGPT. For instance, images with resolutions of 512 $\times$ 512, 256 $\times$ 1024, and 1024 $\times$ 256 can all be encoded into the same number of tokens, making it impossible to infer the original shape without examining the token contents. This ambiguity poses significant challenges for both image perception and generation.
@@
Note that while either the height/width indicators or the <end-of-line> tokens can independently achieve disambiguation, we still use both simultaneously because they have distinct benefits. When generating images, the height/width indicators, generated before any image tokens, pre-determine the shape of the image, aiding Lumina-mGPT in composing the image contents. On the other hand, the <end-of-line> tokens can serve as anchors, offering the 1D tokens sequence with additional explicit spatial information.
@@
A meticulously curated dataset of high-resolution photorealistic image-text pairs is used for FP-SFT. Moreover, the pure-text data from OpenHermess (Teknium, 2023) and the image-to-text data from Mini-Gemini (Li et al., 2024a) are also incorporated during training to prevent catastrophic forgetting. To provide users with a natural way to specify the desired resolution of generated images, we developed the resolution-aware prompt (Figure 2). For each image and its corresponding description, the prompt is structured as follows:
@@
While flexible photorealistic image generation is the primary target of Lumina-mGPT, we find that the resulting model after FP-SFT can be efficiently transferred to a wide spectrum of image understanding and generation tasks. We thus present Omni-SFT, a preliminary exploration toward boosting Lumina-mGPT to a visual generalist. Training data for Omni-SFT consists of the following:
1. Single- and multi-turn language-guided image-editing from MagicBrush (Zhang et al., 2024) and SEED (Ge et al., 2024) (only involving the real-world and multi-turn subsets).
2. Dense prediction tasks, including surface norm estimation from NYUv2 Silberman et al. (2012) and ScanNet Dai et al. (2017), depth estimation from Kitti v2 Cabon et al. (2020) and Sintel Butler et al. (2012), pose estimation from MSCOCO Lin et al. (2014), semantic segmentation data annotated with OneFormer (Jain et al., 2023) on image from Laion (Schuhmann et al., 2022), and grounding data from RefCOCO (Kazemzadeh et al., 2014).
3. In-house spatial-conditional image generation data, with conditions including surface norm, depth, pose, and segmentation.
4. A small fraction of data sampled from those used in the previous FP-SFT process.
As demonstrated in Section 4.2, after Omni-SFT, Lumina-mGPT exhibits a general capability for completing a wide range of tasks other than text-to-image generation, indicating the potential for building a multimodal generalist along this direction.
@@
Though multiple tasks are involved in the SFT process, a unified next-token-prediction loss is used for all of the tasks. As Lumina-mGPT is designed as a chat model, all data are organized into single or multi-turn dialogs, with the loss applied only to the response parts. For all experiments, the AdamW (Loshchilov & Hutter, 2017) optimizer with weight decay = $0.1$ and betas = $(0.9,0.95)$ is used, and the learning rate is set to 2e-5. For stabilizing training, z-loss is applied with weight 1e-5 for both the 7B and the 30B model, and for the 7B model dropout with probability 0.05 is additionally applied. To accommodate the large model volume, PyTorch FSDP (Zhao et al., 2023) is employed with gradient checkpointing. To increase training throughput, all data are pre-tokenized before training and are clustered according to the number of tokens, ensuring that each global batch is composed of data with similar lengths.
3.2.3 Inference for Lumina-mGPT
Classfier-Free Guidance
Classifier-Free Guidance (CFG) (Ho & Salimans, 2022; Sanchez et al., 2023) is originally proposed to enhance the quality and text alignment of generated samples in text-to-image diffusion models. We incorporate this technique into autoregressive models during inference. When generating an image token, the CFG-processed logits $l_{\text{cfg}}$ are formulated as $l_{\text{cfg}}=l+s(l-l^{\prime})$, where $l$ represents the original logits conditioned on the complete context; $l^{\prime}$ represents the context-independent logits, which are conditioned solely on the tokens following the <start-of-image> token of the currently generating image, and are independent of any prior context; $s$ denotes the guidance scale of Classifier-Free Guidance. To make CFG work, during training, the context before <start-of-image> is randomly dropped by a probability of 10%. In practice, KV cache can be used for accelerating the computation of both $l$ and $l^{\prime}$. As shown in Section4.4, CFG has a significant impact on the generation performance of Lumina-mGPT.
Different Decoding Hyperparameters for Image and Text
During inference, the sampling strategy of autoregressive models involves numerous hyperparameters that significantly influence the sampling results. We find that the optimal decoding hyperparameters differ greatly between text decoding and discrete image code decoding. For example, the top-k=5 setting performs well in generating text. However, when generating images, the value of top-k should be much larger (e.g. 2000) to avoid repetitive and meaningless patterns. Therefore, we implement a status-aware control mechanism for inference. Specifically, a set of default hyperparameters is used for text decoding; once a <start-of-image> token is generated, the hyperparameters switch to those optimized for image generation. After the <end-of-image> token is generated, the parameters revert to the initial settings.
Figure 3: Qualitative comparison with LlamaGen and Parti. Lumina-mGPT can generate more photorealistic images with high aesthetics and fine-grained details.
@@ 4 Experiments
4.1 Fundamental Photorealistic Text-to-Image Generation
We first demonstrate the fundamental text-to-image generation capabilities of Lumina-mGPT with FP-SFT. As shown in Figure 1, Lumina-mGPT can generate photorealistic images in a variety of resolutions, achieving the first native 1K autoregressive generation without model cascaden (Yu et al., 2022; Saharia et al., 2022; Chang et al., 2023; Pernias et al., 2024), a common technique in text-to-mage generation. These generated images exhibit strong semantic coherence with intricate visual details, despite being finetuned on limited computational resources and text-image pairs.
Compaison with SoTA AR-based Approaches
We compare the text-to-image synthesis ability of Lumina-mGPT with LlamaGen (Sun et al., 2024) and Parti (Yu et al., 2022). LlamaGen beats state-of-the-art diffusion models on ImageNet FID score. Compared with LlamaGen, Lumina-mGPT can achieve better visual quality as shown in Figure 3 in text-to-image generation. Note that Lumina-mGPT only requires 10M image-text pairs while LlamaGen is trained over 50M low-quality image-text pairs accompanied by 10M in-house aesthetic image-text pairs. Compared to Parti, an AR text-to-image model with 20 billion parameters, Lumina-mGPT also demonstrates better visual quality and aesthetics. However, due to significant differences in computational costs and training datasets, Lumina-mGPT demonstrates inferior text instruction following ability compared to Parti. In addition, neither LlamaGen nor Parti supports the end-to-end generation of 1K resolution images with arbitrary aspect ratios, as achieved by Lumina-mGPT. LlamaGen only supports a fixed resolution of 512 $\times$ 512, while Parti generates 1024 $\times$ 1024 images using an additional super-resolution upsampler.
@@ 4.2 Omnipotent Task Unification with Lumina-mGPT
By performing Omni-SFT over FP-SFT, Lumina-mGPT can perform a multitude of capabilities, which can be categorized into text-only multi-turn dialog, visual multi-turn dialog, multi-turn image editing, dense labeling, and spatial-conditional image synthesis. To intuitively illustrate these capabilities, we qualitatively visualize how various types of downstream tasks can be seamlessly integrated into Lumina-mGPT in Figure 12 to 16.
@@
First, Lumina-mGPT supports general text-only tasks in LLMs, such as solving math problems, coding, and commonsense reasoning, thanks to the extensive pertaining in Chameleon and our multi-task finetuning to mitigate catastrophe forgetting. For instance, Lumina-mGPT correctly answers ‘‘Which is bigger? 9.9 or 9.11’’, as shown in Figure 12, a question that has confused almost all existing LLMs including GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023).
@@
As demonstrated in Figure 13, Lumina-mGPT is also capable of handling various vision-language tasks including image captioning, visual question answering, and general multi-turn multi-image dialog. As a visual generalist, it is important to incorporate classic visual recognition tasks. Using natural language as a unified interface, Lumina-mGPT can perform multiple computer vision tasks including image segmentation, pose estimation, depth estimation, surface normal estimation, and referring object detection. See Figure 14 and 15 for examples.
@@
Beyond photorealistic image generation, it is also vital to add additional control to the text-to-image generation model. Demonstrated in Figure 15 and 16, Lumina-mGPT supports versatile spatial control, including depth maps, segmentation maps, normal maps, and human poses, to guide the generation of target images. Given the above examples, though preliminary, they showcase that Lumina-mGPT can effectively follow diverse instructions, indicating its promising potential as a unification of various challenging tasks in one framework.
4.3 Comparison with Diffusion-based Appraoches
For a long period of time, diffusion models have dominated the field of text-to-image generation compared to autoregressive models. Although LlamaGen claims to beat diffusion models, their results are limited to the ImageNet benchmark and there has been no direct comparison between these two architectures so far. In this section, we aim to provide a detailed comparison of autoregressive and diffusion-based methods trained on these same text-image datasets, focusing on image quality, diversity, text-rendering, and multilingual capabilities. Specifically, we adopt Lumina-mGPT and Lumina-Next-SFT (Zhuo et al., 2024) as representatives of autoregressive and diffusion-based methods, respectively. A direct visual comparison between Lumina-Next-SFT and Lumina-mGPT reveals both the similarities and differences between autoregressive and diffusion-based generative modeling approaches.
On the Similarity between Diffusion- and AR-based Generation Given the same set of text prompts, both diffusion- and AR-based approaches generate photorealistic images with similar aesthetic style and fine-grained details, illustrated in Figure 5. This reveals the fact that both architectures can achieve satisfactory text-to-image generation performance when provided with the same training data, training budget, and comparable model sizes. The AR-based methods display remarkable visual aesthetics on par with their diffusion counterparts, challenging the notion that diffusion models are more effective and promising architecture in generative modeling. This finding also aligns with the platonic representation hypothesis (Huh et al., 2024) that neural networks are converging to learn a shared representation space despite being trained with different architectures and objectives. Therefore, this hypothesis highlights the importance of collecting more high-quality data and optimizing training infrastructure as directions for data and model scaling, to boost the overall model performance that is agnostic to any specific architecture.
On the Differences between Diffusion- and AR-based Generation Despite the visual similarities, there are notable differences between diffusion- and AR-based methods. As shown in Figure 5, Lumina-mGPT exhibits more diversity using different random seeds, while Lumina-Next-SFT generates similar images with identical layouts and textures. This can be partly attributed to the use of high temperature and top-k values in Lumina-mGPT. However, excessive diversity also causes our model to be less stable and more prone to producing visual artifacts, which is discussed in Section 4.7.
In addition to text-to-image generation, Lumina-mGPT supports various vision and language tasks within a unified framework. However, the design of diffusion models limits their compatibility and performance across multiple modalities and tasks. They often require specific architecture designs and additional training for each unseen task (Ke et al., 2024; Xu et al., 2023). In contrast, Lumina-mGPT treats input from all modalities as multimodal token sequences and leverages natural language as the interface to unify diverse tasks with next-token prediction.
4.4 Decoding Configuration Matters
Figure 8: Samples generated by Lumina-mGPT using different decoding configs. Increasing temperature, top-k, and guidance scale all lead to improved visual details.
In autoregressive models, various configuration parameters during Lumina-mGPT’s decoding stage significantly affect sample quality (Holtzman et al., 2020; Radford et al., 2018; 2019). Hyperparameters such as temperature (T), top-k, and classifier-free guidance scale (CFG) have not been extensively investigated in the visual domain. In this section, we explore how these hyperparameters influence the generated image in terms of quality, texture, and style.
Influence of Temperature To evaluate the effect of these decoding parameters, we first set a standard decoding configuration: T=1.0, Top-k=2000, CFG=4.0, which serves as a good-to-use setting. From this baseline, we gradually shift T from 0.7 to 1.0 to generate corresponding images at different temperatures. The results are shown in Figure8. It is evident that when setting the temperature low, visual details diminish and objects tend to be over-smoothed. Conversely, when setting the temperature high, the generated images contain rich visual content but are prone to contain more artifacts.
Influence of Top-k Based on the standard decoding setting, we vary the top-k value, from 50 to 8192, where 8192 is equal to the size of the VQ-VAE codebook used usedin Lumina-mGPT. The results, visualized in Figure8, indicate a similar trend with increasing temperature. When top-k is low, the image content and texture are relatively simple, exhibiting the over-smoothed problem as well. When top-k is set high, the image detail and texture are diverse, making it more aesthetically appealing, while increasing the potential of artifacts.
Influence of Classifier-Free Guidance CFG is extremely important in text-to-image diffusion models (Ho & Salimans, 2022; Sanchez et al., 2023). To verify its effectiveness in autoregressive generation, we change the CFG value from 1.0 to 8.0. As shown in Figure8, with the CFG increasing, the quality of generated images improves, proving the effectiveness of the classifier-free guidance in this context.
4.5 Lumina-mGPT as VQ Codes Refiner
Figure 9: Reconstruction quality comparison between different methods. Lumina-mGPT can be instructed to refine the encoded VQ codes, leading to better reconstructed images.
It is well-known that both VQ-VAE (Van Den Oord et al., 2017; Razavi et al., 2019; Esser et al., 2021) and VAE (Kingma & Welling, 2013) compress images into latent representations at the cost of information loss, which introduces quality degradation at reconstruction, especially for high-frequency details such as edges, hair, and text. This evitable loss upper bounds the quality of image generation for all latent space generative models (Esser et al., 2024; Yu et al., 2022; Sun et al., 2024). However, we surprisingly found that Lumina-mGPT can improve the discrete image tokens encoded by VQ-VAE using a simple ‘‘No edit. <|image|>’’ instruction, then Lumina-mGPT can generate a sequence of refined image tokens that can be decoded into a better image.
We compare the reconstruction quality of VQ-VAE used in Chameleon, Chameleon’s VQ-VAE using VQ codes refined by Lumina-mGPT, and the SDXL VAE (Podell et al., 2023) in Figure 9. The SDXL VAE showcases significantly better reconstruction performance compared to Chameleon’s VQ-VAE. This advantage also explains why diffusion-based methods can beat autoregressive models in text-to-image generation. However, after applying the aforementioned simple trick to refine the VQ codes using Lumina-mGPT, the quality of reconstructed images exhibits remarkable gains in both fine-grained visual details and text rendering, on par with the SDXL VAE. We hypothesize that Lumina-mGPT learns the underlying distribution of real image tokens during large-scale pertaining and high-quality FP-SFT. The Omni-SFT further triggers the emergence of such zero-shot capability after training on various tasks and instructions. We believe this zero-shot capability to bridge the representation gap between the encoder and decoder in VQ-VAE suggests promising directions to improve the current paradigm of two-stage latent generative modeling. For example, we can leverage Lumina-mGPT as a teacher to refine the VQ codes for student training or to design inference-time techniques by manipulating the token logits similarly to classifier-free guidance.

### v3

@@ 2.2.1 Effective Initlization
Large-scale pre-training and scalable model architecture have been widely verified as the golden path to advanced intelligence. As mGPTs like Chameleon are pretrained on large-scale datasets and have developed effective and generalizable representations, they are well-suited to serve as the starting point for flexible photorealistic image generation and beyond. In this work, we initialize from Chameleon with parameters ranging from 7B to 30B, using just 10M high-quality image-text data points.
@@ 2.2.2 Supervised Finetuning for Lumina-mGPT
Existing methods represent images as 1D flattened sequences of 2D discrete image codes. While adequate for fixed resolutions, this approach becomes ambiguous when supporting variable resolutions, as with Lumina-mGPT. For instance, images with resolutions of 512 $\times$ 512, 256 $\times$ 1024, and 1024 $\times$ 256 can all be encoded into the same number of tokens, making it impossible to infer the original shape without examining the token contents. This ambiguity poses significant challenges for both image perception and generation.
@@
Discussions: While either the height/width indicators or the <end-of-line> tokens can independently achieve disambiguation, we use both simultaneously due to their distinct benefits. When generating images, the height/width indicators, generated before any image tokens, pre-determine the shape of the image, aiding Lumina-mGPT in composing the image contents. The <end-of-line> tokens, on the other hand, serve as anchors, offering the 1D token sequence with additional explicit spatial information. Section 3.3 shows empirical analysis over these tokens.
@@
A meticulously curated dataset of high-resolution photorealistic image-text pairs is used for FP-SFT. Moreover, the pure-text data from OpenHermess (OpenHermes2.5) and the image-to-text data from Mini-Gemini (mgm) are also incorporated during training to prevent catastrophic forgetting. To provide users with a natural way to specify the desired resolution, we develop a resolution-aware prompt (Figure 2). For each image and its corresponding description, the prompt is structured as follows:
@@
While flexible photorealistic image generation is the primary target of Lumina-mGPT, we find that the resulting model after FP-SFT can be efficiently transferred to a wide spectrum of image understanding and generation tasks. We thus present Omni-SFT, a preliminary exploration toward boosting Lumina-mGPT to a visual generalist. Training tasks and data for Omni-SFT consists of the following:
1. Single- and multi-turn language-guided image-editing with data from MagicBrush (zhang2024magicbrush) and SEED (SeedDataEdit) (only involving the real-world and multi-turn subsets).
2. Dense prediction tasks, including surface norm estimation from NYUv2 (silberman2012indoor) and ScanNet (dai2017scannet), depth estimation from Kitti v2 (cabon2020virtual) and Sintel (butler2012naturalistic), pose estimation from MSCOCO (lin2014microsoft), semantic segmentation data annotated with OneFormer (jain2023oneformer) on image from Laion (schuhmann2022laion), and grounding data from RefCOCO (kazemzadeh2014referitgame).
3. In-house spatial-conditional image generation following ControlNet (controlnet), with conditions including surface norm, depth, pose, and segmentation.
4. Text-conditional multiview generation using an internal dataset consisting of  100k high-quality samples with rendered $384^{2}$ images from 16 viewpoints uniformly distributed in azimuth angles.
5. A small fraction of data sampled from those used in the previous FP-SFT process, including both text modeling and text-to-image generation to maintain its learned capabilities.
We tokenize all text and images into discrete tokens and formulate these tasks as a unified next-token prediction objective. Notably, we also incorporate tasks multiview generation, which requires generating a sequence of image frames, as a preliminary for video generation. As demonstrated in Section 3.2, after Omni-SFT, Lumina-mGPT exhibits a general capability for completing a wide range of tasks other than text-to-image generation, indicating the potential for building a multimodal generalist along this direction.
Figure 3: Qualitative comparison on text-to-image generation. Prompts: (1) A raccoon wearing formal clothes, wearing a tophap and holding a cane. The raccoon is holding a garbage bag. Oil painting in the style of Rembrandt. (2) A teddy bear wearing a motorcycle helmet and cape is standing in front of Loch Awe with Kilchurn Castle behind him, dslr photo. (3) A tornado made of sharks crashing into a skyscraper. Painting in the style of Hokusai. (4) A photo of an Athenian vase with a painting of pandas playing basketball in the style of Egyptian hieroglyphics.
@@
Though multiple tasks are involved in the SFT process, a unified next-token-prediction loss is used for all of the tasks. As Lumina-mGPT is designed as a chat model, all data are organized into single or multi-turn dialogs, with the loss applied only to the response parts. For all experiments, the AdamW (adamw) optimizer with weight decay = $0.1$ and betas = $(0.9,0.95)$ is used, and the learning rate is set to 2e-5. Inspired by the classifier-free guidance in diffusion models (cfg), we randomly drop the context by a probability of 10% during training, as detailed in Appendix F. To accommodate the large model volume, PyTorch FSDP (fsdp) is employed with gradient checkpointing. To increase training throughput, all data are pre-tokenized before training and are clustered according to the number of tokens, ensuring that each global batch is composed of data with similar lengths.
To stabilize training, we applied z-loss (palm) with a weight of $1\times 10^{-5}$ for both the 7B and 30B models. Additionally, dropout with a probability of 0.05 was applied specifically to the 7B model. Initially, we underestimated the significance of z-loss, as it is rarely mentioned in the literature on training multimodal large language models. However, our experiments revealed that omitting this term led to a surge in the magnitude of logits, resulting in diverging loss curves. Incorporating z-loss proved to be pivotal in enhancing training stability. Furthermore, we observed that applying z-loss significantly reduced the magnitude of logits, which in turn lowered the optimal temperature for inference-time image generation.
Table 2: Quantitative comparison on text-to-image benchmarks.
Methods GenEval $\uparrow$ DPG $\uparrow$ T2I-CompBench $\uparrow$
Two Obj. Counting Color Attri. Overall Entity Relation Attribute Overall Color Shape Texture
Diffusion Models
SDv1.5 (sd2_1) - - - 0.40 74.23 73.49 75.39 63.18 0.3730 0.3646 0.4219
Lumina-Next (lumina-next) 0.49 0.38 0.15 0.46 83.78 89.78 82.67 75.66 0.5088 0.3386 0.4239
SDv2.1 (sd2_1) 0.51 0.44 0.50 0.47 - - - 68.09 0.5694 0.4495 0.4982
SDXL (sdxl) 0.74 0.39 0.23 0.55 82.43 86.76 80.91 74.65 0.6369 0.5408 0.5637
SD3-medium (sd3) 0.74 0.63 0.36 0.62 91.01 80.70 88.83 84.08 - - -
DALL-E3 (dalle3) 0.87 0.47 0.45 0.67 89.61 90.58 88.39 83.50 0.8110 0.6750 0.8070
AutoRegressive Models
LlamaGen (llamagen) 0.34 0.21 0.04 0.32 - - - 65.16 - - -
Chameleon (chameleon) - - - 0.39 - - - - - - -
Lumina-mGPT (Ours) 0.77 0.27 0.32 0.56 86.60 91.29 84.61 79.68 0.6371 0.4727 0.6034
@@ 3 Experiments
3.1 Photorealistic Text-to-Image Generation
We first demonstrate the fundamental text-to-image generation capabilities achieved after FP-SFT. As shown in Figure 6, Lumina-mGPT can generate photorealistic images in a variety of resolutions, achieving the first native 1K autoregressive generation without cascaded models (parti; imagen; muse; wurstchen). Though finetuned on limited computational resources and text-image pairs, the results exhibit strong semantic coherence with intricate visual details
Qualitative Comparison
Compared with LlamaGen (llamagen) and Parti (parti), Lumina-mGPT achieves better visual quality and aesthetics (Fig. 3). Besides, LlamaGen and Parti do not support end-to-end generation at high resolutions (e.g., 1k) or at arbitrary aspect ratios, whereas it is supported by Lumina-mGPT. Specifically, LlamaGen only supports a fixed resolution of 512 $\times$ 512, while Parti generates 1024 $\times$ 1024 images using an additional super-resolution upsampler. Beyond AR-based approaches, we also provide a side-by-side comparison with diffusion-based counterparts by training on the same dataset, as detailed in Appendix G.
Quantitative Comparison
We evaluate the performance of Lumina-mGPT on widely used text-to-image benchmarks T2I-CompBench (huang2023t2icompbench), GenEval (GenEval), and DPG-Bench (hu2024ella). The results are presented in Tab. 2.
Notably, Lumina-mGPT shows significant improvement over Chameleon (chameleon) and surpasses Lumina-Next (lumina-next), a state-of-the-art diffusion transformer trained on the same text-to-image dataset as Lumina-mGPT. Moreover, Lumina-mGPT consistently outperforms SDv1.5 and SDv2.1 (sd2_1), becoming the first autoregressive model with competitive generation quality to SDXL (sdxl). Nevertheless, a clear gap remains towards SOTA diffusion models like SD3 (sd3) and DALL-E3 (dalle3), showing the room for further progress along this direction.
@@ 3.1 Photorealistic Text-to-Image Generation
Decoding Configuration Matters
We observe the significant difference between the optimal hyperparameters for text and image generation, which inspires us to introduce a status-aware control mechanism switching settings based on the type of content being generated. We then explore how different inference configurations, such as temperature, top-k, and classifier-free guidance scale, affect the quality of generated images. For example, lower temperatures and top-k values often result in over-smoothed images, while higher values enhance detail but may introduce artifacts. Due to space limits, we defer the details to Appendix F.
@@
Table 3: Evaluation on image understanding benchmarks.
Method MMBench MME-p MME-r
Chameleon 19.80 153.10 49.60
Lumina-mGPT 32.20 976.85 290.36
Method SEEDBench-I MMMU-val POPE
Chameleon 30.50 22.40 19.40
Lumina-mGPT 50.93 27.11 70.43
@@ 3.2 Omnipotent Task Unification with Lumina-mGPT
Beyond text-to-image generation, Lumina-mGPT also possesses the capability of text-only dialog and image understanding. After Omni-SFT, the spectrum of capabilities is further extended, covering multi-turn image editing, dense labeling, spatial-conditional image synthesis, and multiview generation. Such unification embodies the inherent benefits and potential of the decoder-only autoregressive architecture for unified modeling. Notably, the high-quality image generation ability achieved by Lumina-mGPT serves as the bedrock for such unification, without which the relevant explorations would be impossible.
To intuitively illustrate these capabilities, we provide visualizations in Figure 7 to 11 in the Appendix.
@@
First, Lumina-mGPT works effectively as a normal LLM to handle general text-only tasks (Figure 7), such as math problem solving, coding, and commonsense reasoning, thanks to the extensive pertaining in Chameleon and our involvement of text-only data to mitigate catastrophic forgetting.
Shown in Figure 8, Lumina-mGPT can also handle various visual understanding tasks including captioning, visual question answering, and general multi-turn multi-image dialog. Akin to image generation, flexible image aspect ratios are also supported in image understanding. Quantitative results on visual understanding benchmarks (mmbench; mmmu; pope; seedbench; mme) are presented in Table 3, illustrating that Lumina-mGPT significantly outperforms the original Chameleon model.
@@
As a visual generalist, Lumina-mGPT incorporates classic visual recognition tasks. Using natural language as a unified interface, Lumina-mGPT can perform multiple high-level computer vision tasks including image segmentation, pose estimation, depth estimation, surface normal estimation, and referring object detection. See Figure 9 and 10 for examples.
Lumina-mGPT also supports image generation with versatile spatial conditions, such as depth maps, segmentation maps, normal maps, and human poses, as demonstrated in Figure 10 and 11. Beyond single-image generation, Lumina-mGPT can produce a sequence of consistent images from multiview, as shown in Figure 11. This task marks a preliminary step toward video generation by rendering the multiview images into a coherent video.
@@
Given the above examples, though preliminary, they showcase that Lumina-mGPT can effectively follow diverse instructions, highlighting its promising potential as a unification of various challenging tasks in one framework.
Figure 5: Visualization of averaged attention logits from the last image token. Added indicator tokens are allocated with a large proportion of attention scores.